OpenStudy (anonymous):

0=-cosx^2-sinx+sinx^2 stuck on how to solve this...

OpenStudy (anonymous):

scratch that. lets stick witht he trig functions

OpenStudy (anonymous):

\[0=-\cos x^2-\sin x+\sin x^2\]\[\cos x^2 = \sin x^2 - \sin x\] http://www.sosmath.com/trig/Trig5/trig5/trig5.html \[1-\sin x ^2=\sin x^2 - sin x\]

OpenStudy (anonymous):

\[0=2\sin x^2 - \sin x-1\]

OpenStudy (anonymous):

\[0=(2\sin x +1)(\sin x - 1)\]

OpenStudy (anonymous):

now we have to think of the zeros.

OpenStudy (anonymous):

are you familiar with the unit circle

OpenStudy (anonymous):

yes! thanks so much! appreciate your help!

OpenStudy (anonymous):

Do you still need help? At the very least I'll check your answer if you post it

OpenStudy (anonymous):

ok. im working on it.

OpenStudy (anonymous):

there should be 3 answers

OpenStudy (anonymous):

ok im confused how you got from \[2sinx ^{2}-\sin x-1\] to 0=(2sinx+1)(sinx−1)

OpenStudy (anonymous):

First: I used trig properties, see the website I posted, to get the equation in the same term (sin x) Second, I set the the equation to 0 so I could factor. I could have also used the quadratic. Do you see that?

OpenStudy (anonymous):

if x = sin x, then \[0=2x^2-x-1\]

OpenStudy (anonymous):

solve that then you can get the zeros

OpenStudy (anonymous):

OH. mygoodness. i didn't see that. ok. the quadratic makes so much more sense now

OpenStudy (anonymous):

Be careful because the quadratic just helps get it into the factored form, but because it is a trig function you need to think about the unit circle. I'll show the quadratic steps

OpenStudy (anonymous):

yeah. with you showing the quadratic form i understand how its factored. so now i set each factor to zero and solve. sinx=-1/2 sinx=1 so x is 7pi/6 and pi/2 right?

OpenStudy (anonymous):

oh and 11pi/6

OpenStudy (anonymous):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]\[x=\frac{1 \pm \sqrt{1-4*2*-1}}{2*2}\]\[x=\frac{1 \pm 3}{4}\] yes good job. Do you understand what I was saying about the quadratic just being a tool in this case and not the zeros?

OpenStudy (anonymous):

yeah. showing how it could be factored but leaving it in the trig form. Thank you so much!

OpenStudy (anonymous):

\[x=\frac{1 \pm 3}{4}\]\[x=\frac{-1}{2},1\]\[(\sin x + \frac{1}{2})(\sin x - 1)\]so where you would have an x you re-replace it with sinx

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

I'd print out that webpage I posted. That site has helped me so much with trig problems

OpenStudy (anonymous):

i will! 'course, those only help if i can see the connections. thanks again!

OpenStudy (anonymous):

It helps me to think about the problem backwards. What am i solving for? What do I need to do to get that alone? that's what I think about

OpenStudy (anonymous):

ok. that will definitely help.